OpenStudy (anonymous):

Find the monthly payment needed to have a sinking fund accumulate a future value of $18,000. The yearly interest rate is 10.8% and the number of payments is 21. Interest is compounded monthly. In your calculations, do not round until the final answer. Then, round the monthly payment to the nearest cent.

OpenStudy (paki):

with payments at the beginning of each month... FV = PMT * (1.009)(1.009^21-1) / .009 PMT = (18000*.009 /1.009)/(1.009^21-1) = $775.55 <-------- if payments are at the end of each month, $782.53 <---------

OpenStudy (anonymous):

ok so 782.53 makes it the final answer? @paki

OpenStudy (paki):

yeah... i have written it...

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

btw, do you know how I can insert the formulas into the TI-84 calculator? For ex. A=R[(1+r/m)mt-1/r/m] ?

OpenStudy (paki):

pleasure